OpenStudy (anonymous):

Quick Help! So I have to find the equation of the line that passes through (3,2) and (5,6) I figured out that the slope is 2 and now need to find B in the y=mx+b equation. Help?

OpenStudy (anonymous):

two-point form p1(3,2) p2(5,6) formula (y-y1)(x2-x1)=(y2-y1)(x-x1) (y-2)(5-3)=(6-1)(x-3) (y-2)(2)=(5)(x-3) 2y-4=5x-15 -5x+2y=-15+(-4) -5x+2y=-19 <- standard form 2y=5x-19<- not yet simplified 2y/2=5x/2-19/2 y=5/2 x - 19/2 <-simplified form m=5/2 and b= 19/2

OpenStudy (anonymous):

hope this helps